# Quantised helicity in optical media

### v1

@@ Abstract
Optical helicity quantifies the handedness of light, and plays a central role in the description of interactions between light and chiral matter. In free space, it is related to the duality symmetry of the electromagnetic field, a continuous symmetry encapsulating the invariance of Maxwell’s equations under the interchange of electric and magnetic fields. However, in materials the situation is not so straightforward, as the free space transformation must be extended to encompass mixing of both the $\mathbf{E}$/$\mathbf{H}$ and $\mathbf{D}$/$\mathbf{B}$ field pairs. The simultaneous direct interchange of $\mathbf{E}$/$\mathbf{H}$ and of $\mathbf{D}$/$\mathbf{B}$ is incompatible with the presence of linear constitutive relations. In this work, we extend the duality transform in a way that resolves this incompatibility, and use this to define the optical helicity in a general medium, which may be dispersive, lossy, chiral or nonreciprocal. We find that the helicity density must contain an explicit contribution associated with the polarisation and magnetisation of the matter, and we show that the form of this matter contribution is independent of the details of the medium. We also show that the in-medium helicity can be naturally expressed in terms of the elementary quantised excitations of the system.
@@
×$\boldsymbol{\nabla}\times\mathbf{E}$ $=-\partial_{t}\mathbf{B},$$\boldsymbol{\nabla}\times\mathbf{H}$ $=\partial_{t}\mathbf{D}.$ (1)
It was Heaviside and Larmor [9, 10] who pointed out that these equations possess an electric-magnetic symmetry in that they are unchanged if we perform the rotations
@@
where $\theta$ is a pseudoscalar, and we work in natural units in which $\varepsilon_{0}$, $\mu_{0}$, $c$ and $\hbar$ are all unity. In free space we can associate this symmetry with a locally conserved helicity density
@@
where $\mathbf{A}$ is the magnetic vector potential, defined by $\mathbf{B}=\boldsymbol{\nabla}\times\mathbf{A}$, and $\mathbf{C}$ is the electric vector potential, defined by $\mathbf{D}=-\boldsymbol{\nabla}\times\mathbf{C}$ [11, 12]. The flux or flow of the helicity is simply the spin density [13, 14]. Indeed, in free space the Heaviside-Larmor transformation (Quantised helicity in optical media) may be interpreted as a rotation of the fields for each plane wave through an angle $\theta$ about its direction of propagation. As required the total helicity, which is the volume integral of $h_{\text{free}}$, acts as the generator of the transformation.
An important consequence of the Heaviside-Larmor symmetry is that all physical quantities must also be invariant under it [15]. This means, for example, that the energy density $(\mathbf{E}\cdot\mathbf{D}+\mathbf{B}\cdot\mathbf{H})/2$, and the kinetic and canonical momentum densities $\mathbf{E}\times\mathbf{H}/c^{2}$ and $\mathbf{D}\times\mathbf{B}$ [16], are satisfactory, but that $\varepsilon_{0}\mathbf{E}\times\mathbf{B}$ is not. This condition applies also to optical forces on a medium, and we note that the the Einstein-Laub force density [17]
@@
also satisfies the symmetry.
@@
which is not the form required by the Heaviside-Larmor symmetry. This problem is depicted for both $\mathbf{D^{\prime}}$ and $\mathbf{B}^{\prime}$ in Figure 1. The transformations (Quantised helicity in optical media) and (Quantised helicity in optical media) are compatible only when $\varepsilon/\mu=1$. In this case, we say that the medium is dual-symmetric.
One way in which the incompatibility of (Quantised helicity in optical media), (Quantised helicity in optical media) and the constitutive relations has been addressed previously is by defining a duality transform in media which explicitly includes the permittivity and permeability [18, 19, 20, 21]. That is, to use the transformation
@@
in place of (Quantised helicity in optical media). This transformation is a symmetry of Maxwell’s equations (1) under the condition that $\boldsymbol{\nabla}\sqrt{\varepsilon/\mu}=\mathbf{0}$ throughout space. This approach leads to an optical helicity density similar to the free space (3), with the addition of appropriate factors of $\varepsilon$ and $\mu$ [20]. In chiral media, the situation is more complicated, and the helicity density derived from the transformation above includes extra contributions proportional to the energy density in the medium [21].
Figure 1: Illustration comparing two ways of defining the duality transform in a medium. (a) The duality transform defined in equation (Quantised helicity in optical media) implies a simple rotation between the $\mathbf{D}$ and $\mathbf{B}$ fields; (b) The duality transforms defined in equations (Quantised helicity in optical media) and (Quantised helicity in optical media) are equivalent to a rotation and scaling, and alter both the relative lengths of $\mathbf{D}^{\prime}$ and $\mathbf{B^{\prime}}$ and the angle between them. It is clear that these transformations are both incompatible with those in equation (Quantised helicity in optical media).
The difference between the transformations (Quantised helicity in optical media) and (Quantised helicity in optical media) can be emphasised by noting that (Quantised helicity in optical media) corresponds to a rotation between $\mathbf{D}$ and $\mathbf{B}$, while (Quantised helicity in optical media) corresponds to a rotation and scaling, with a scale factor which depends on both $\varepsilon$ and $\mu$. Dispersion can be accommodated for in the transformation (Quantised helicity in optical media) [20]. From this point of view, treating dispersive media amounts to applying a different scaling for the different frequency components of the fields.
This approach works for linear, lossless materials, but there are features of the transformation (Quantised helicity in optical media) that make it very difficult to extend to more general cases. It is unclear how to account for frequency dependent absorption. Nonlinear responses are excluded from the outset, as the definition of the transformation (Quantised helicity in optical media) assumes that the field pairs $\mathbf{D}/\mathbf{B}$ and $\mathbf{E}/\mathbf{H}$ are connected by linear constitutive relations.
Furthermore, we may note that the transformations (Quantised helicity in optical media) imply the following transformation of $\mathbf{P}$ and $\mathbf{M}$,
@@
We therefore see that, while the transformations (Quantised helicity in optical media) may be a symmetry of Maxwell’s equations (1) even when $\varepsilon\neq\mu$, they then imply transformations of $\mathbf{P}$ and $\mathbf{M}$ that depend explicitly on the fields, rather than the simple rotation implied by (Quantised helicity in optical media).
All of these issues stem from the fact that (Quantised helicity in optical media) explicitly includes the constitutive parameters $\varepsilon$ and $\mu$ in the transformation. The approach of this Letter is to eschew (Quantised helicity in optical media), and to instead consider
@@
as the duality transformation from which to derive the optical helicity inside a general medium. The key to this approach is that the fields $\mathbf{P}$ and $\mathbf{M}$ are not derived from the imposition of constitutive relations, but are explicitly included in the Lagrangian through an effective model of the material’s internal degrees of freedom (as in [22, 23]). The explicit inclusion of the material degrees of freedom in the Lagrangian leads to the natural emergence of a matter component in the medium helicity.
In order to enforce both of the transformations in Eq. (Quantised helicity in optical media) and, thereby, generate the Heaviside-Larmor transformation (Quantised helicity in optical media), we need to extend the form of the helicity and of its flux to include both field and matter components. This suggests that the helicity density should take the form
$h=\frac{1}{2}\left(\mathbf{A}\cdot\mathbf{B}-\mathbf{C}\cdot\mathbf{D}\right)+\boldsymbol{\Pi}^{P}\cdot\mathbf{M}-\boldsymbol{\Pi}^{M}\cdot\mathbf{P},\$ (9)
where $\boldsymbol{\Pi}^{P}$ and $\boldsymbol{\Pi}^{M}$ are the canonical momenta conjugate to $\mathbf{P}$ and $\mathbf{M}$. It is straightforward to confirm that the volume integral of this density generates the full Heaviside-Larmor transformation. If we promote our fields to quantum mechanical operators, then we find
@@
to first order in $\theta$. The infinitesimal transformations of $\mathbf{D}$ and $\mathbf{B}$ follow directly. Finally, the flux density of our generalised helicity density is simply
@@
which we recognise as the density of the electromagnetic spin.
For situations in which the helicity is locally conserved, the helicity density and this flux satisfy the local conservation law
@@
as may be derived directly from Noether’s theorem. The details of this are given in the Supplemental Material [24].
@@
In a magnetodielectric with a single electric and single magnetic resonance, the polarisation and magnetisation fields satisfy the equations of motion
@@
where $\omega_{E}$ and $\omega_{M}$ are the resonance frequencies, and $\alpha$ and $\beta$ are constants. With time-dependence $e^{-i\omega t}$, it is clear that these equations imply the constitutive relations $\mathbf{D}=\varepsilon\mathbf{E}$ and $\mathbf{B}=\mu\mathbf{H}$, with
@@
The equations of motion (16) are only invariant under the duality transform (Quantised helicity in optical media) if $\alpha=\beta$ and $\omega_{E}=\omega_{M}$. This means that $\varepsilon(\omega)=\mu(\omega)$ at all frequencies, and therefore that the electric and magnetic constituents of the medium emit in phase. The equations (16), along with Maxwell’s equations, may be obtained from the following Lagrangian density,
@@
$+\frac{1}{2\alpha^{2}}\big{[}(\partial_{t}\mathbf{P})^{2}-\omega_{E}^{2}\mathbf{P}^{2}\big{]}+\frac{1}{2\beta^{2}}\big{[}(\partial_{t}\mathbf{M})^{2}-(\omega_{M}^{2}+\beta^{2})\mathbf{M}^{2}\big{]}.$
We take this Lagrangian as a starting point from which to derive an expression for the optical helicity. If we set $\alpha=\beta$ and $\omega_{E}=\omega_{M}=\omega_{0}$ then we can apply Noether’s theorem to recover the helicity directly [24]. Our Lagrangian provides the precise forms of the canonical momenta corresponding to the polarisation and magnetisation:
@@
$\mathbf{\Pi}^{M}\equiv\frac{\partial\mathcal{L}}{\partial\dot{\mathbf{M}}}=\partial_{t}\mathbf{M}/\beta^{2},$ (20)
and these may therefore be interpreted as bound polarisation and magnetisation currents for the Lagrangian (18). In the case of more general Lagrangians, this interpretation is not always possible. However, the expression for the helicity, when written in terms of the fields and their conjugate momenta, will still be given by equation (9).
We turn next to a microscopic quantum description of the helicity. In free space, the total helicity can be expressed as the difference between the number of left- and right-handed circularly polarised photons [25],
^$\widehat{\mathcal{H}}_{\text{free}}$ $=\int d^{3}r\,\hat{h}_{\text{free}}=\int d^{3}k\sum_{\lambda=\pm}\lambda\;\hat{a}^{\dagger}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k}),$ (21)
where $\hat{a}_{\lambda}^{\dagger}(\mathbf{k})$ and $\hat{a}_{\lambda}(\mathbf{k})$ are the creation/annihilation operators for a photon of wavevector $\mathbf{k}$, and $\lambda=\pm 1$ labels the two circular polarisations. In a dual-symmetric medium, the helicity operator defined by equation (9) admits a similar interpretation: it is equal to the difference between the number of left- and right-handed polaritons in the system.
To see this, we diagonalise the system and examine the polariton modes. From the Lagrangian (18), we obtain the Hamiltonian

### v2

@@ Abstract
We present a new approach to the definition of optical helicity in a medium. Our approach resolves the problem that duality transformations which simultaneously combine $\mathbf{E}$ with $\mathbf{H}$ and $\mathbf{D}$ with $\mathbf{B}$ are incompatible with linear constitutive relations. We find that the helicity density in a medium, as the conserved quantity associated with duality transforms, must contain an explicit contribution associated with the polarisation and magnetisation of the matter, and that it can be expressed naturally in terms of the elementary polarised excitations of the system. In media for which the helicity is conserved, each circular excitation carries a well-defined helicity. However, in a medium for which the helicity is not conserved, we find that the time-varying helicity can be viewed in terms of oscillations between different helicity eigenstates, analogous to neutrino oscillations. Here we explicitly study the helicity in homogeneous and lossless media but we believe that, differently to other choices, this helicity is readily generalisable to media that may be inhomogeneous, lossy, chiral or nonreciprocal.
I Introduction
@@
$\boldsymbol{\nabla}\times\mathbf{E}$ $=-\partial_{t}\mathbf{B},$ $\boldsymbol{\nabla}\times\mathbf{H}$ $=\partial_{t}\mathbf{D}.$ (1)
It was Heaviside and Larmor [9, 10] who first pointed out that these equations possess an electric-magnetic symmetry in that they are unchanged if we perform the rotations
@@
where $\theta$ is a pseudoscalar, and we work in natural units in which $\varepsilon_{0}$, $\mu_{0}$, $c$ and $\hbar$ are all unity. In free space we can associate this symmetry with a locally conserved helicity density [11]
@@
where $\mathbf{A}$ is the magnetic vector potential, defined by $\mathbf{B}=\boldsymbol{\nabla}\times\mathbf{A}$, and $\mathbf{C}$ is the electric vector potential, defined by $\mathbf{D}=-\boldsymbol{\nabla}\times\mathbf{C}$ [12, 13]. The flux or flow of the helicity is simply the spin density [14, 11]. Indeed, in free space the Heaviside-Larmor transformation (I) may be interpreted as a rotation of the fields for each plane wave through an angle $\theta$ about its direction of propagation. As required the total helicity, which is the volume integral of $h_{\text{free}}$, acts as the generator of this transformation.
An important consequence of the Heaviside-Larmor symmetry is that all physical quantities must also be invariant under it [15]. This means, for example, that the energy density $(\mathbf{E}\cdot\mathbf{D}+\mathbf{B}\cdot\mathbf{H})/2$, and the kinetic and canonical momentum densities $\mathbf{E}\times\mathbf{H}$ and $\mathbf{D}\times\mathbf{B}$ [16], are satisfactory, but that $\mathbf{E}\times\mathbf{B}$ is not, and so lacks the physical significance of the other cross-products. This condition also applies to optical forces on a medium, and we note that the the Einstein-Laub force density [17, 18]
@@
also satisfies the symmetry. Here, $\mathbf{P}$ and $\mathbf{M}$ are the polarisation and magnetisation densities within the medium.
@@
which is not of the form required by the Heaviside-Larmor symmetry. This problem is depicted for both $\mathbf{D^{\prime}}$ and $\mathbf{B}^{\prime}$ in Figure 1. The situation is worse in chiral media, for which $\mathbf{D}$ depends on both $\mathbf{E}$ and $\mathbf{H}$, while $\mathbf{B}$ is a function of both $\mathbf{H}$ and $\mathbf{E}$ [19, 20]. The transformations (I) and (I) are compatible only when $\varepsilon/\mu=1$. In this case, we say that the medium is dual-symmetric.
II Defining a Duality Transform
One way in which the incompatibility of (I), (I) and the constitutive relations has been addressed previously is to define a duality transform in media which explicitly includes the permittivity and permeability [21, 22, 23, 24]. That is, to use the transformation
@@ II Defining a Duality Transform
in place of (I). This transformation is a symmetry of Maxwell’s equations (1) under the condition that $\boldsymbol{\nabla}\sqrt{\varepsilon/\mu}=\mathbf{0}$ throughout space. This approach leads to an optical helicity density similar to the free space expression (3), with the addition of appropriate factors of $\varepsilon$ and $\mu$ [23]. In chiral media, however, the situation is more complicated, and the helicity density derived from this transformation above includes extra contributions proportional to the energy density in the medium [24].
Figure 1: Illustration comparing two ways of defining the duality transform in a medium. (a) The duality transform defined in equation (I) implies a simple rotation between the $\mathbf{D}$ and $\mathbf{B}$ fields; (b) The duality transforms defined in equations (I) and (II) are equivalent to a rotation and scaling, and alter both the relative lengths of $\mathbf{D}^{\prime}$ and $\mathbf{B^{\prime}}$ and the angle between them. It is clear that the transformations (I) and (II) are both incompatible with those in equation (I).
The difference between the transformations (I) and (II) can be emphasised by noting that (I) corresponds to a rotation between $\mathbf{D}$ and $\mathbf{B}$, while (II) corresponds to a rotation and scaling, with a scale factor which depends on both $\varepsilon$ and $\mu$. Dispersion in the absence of absorption can be accommodated for in the transformation (II) [23]. From this point of view, however, treating dispersive media amounts to applying a different scaling for the different frequency components of the fields.
This approach works for linear, lossless materials. However, there are features of the transformation (II) that make it very difficult to extend to more general cases. It is unclear how to account for frequency-dependent absorption, which prohibits the treatment of realistic media that obey the Kramers-Kronig relations. Nonlinear responses are excluded from the outset, as the definition of the transformation (II) assumes that the field pairs $\mathbf{D}/\mathbf{B}$ and $\mathbf{E}/\mathbf{H}$ are connected by linear constitutive relations.
Furthermore, we may note that the transformations (II) imply the following transformation of the polarisation $\mathbf{P}$, and magnetisation $\mathbf{M}$,
@@
We see that, while the transformations (II) may be a symmetry of Maxwell’s equations (1) even when $\varepsilon\neq\mu$, they then imply transformations of $\mathbf{P}$ and $\mathbf{M}$ that depend explicitly on the fields, rather than the simple rotation implied by (I). [^1] There are some cases in which the transformation (II) can be viewed as providing a scaled rotation between $\mathbf{P}$ and $\mathbf{M}$. If we have $\mathbf{D}=\varepsilon\mathbf{E}$ and $\mathbf{B}=\mu\mathbf{H}$, then we can write $\mathbf{P}=\chi_{e}\mathbf{E}$ and $\mathbf{M}=\chi_{m}\mathbf{H}$ with $\varepsilon=1+\chi_{e}$ and $\mu=1+\chi_{m}$. Using these relationships in (II), we obtain $\mathbf{P}^{\prime}=\cos{\theta}\mathbf{P}+\sin\theta\sqrt{\frac{\varepsilon}{\mu}}\frac{\chi_{e}}{\chi_{m}}\mathbf{M},$ $\mathbf{M}^{\prime}=\cos{\theta}\mathbf{M}-\sin{\theta}\sqrt{\frac{\mu}{\varepsilon}}\frac{\chi_{m}}{\chi_{e}}\mathbf{P},$ which, for a plane wave propagating through the medium, would effect a physical rotation of the $\mathbf{P}$ and $\mathbf{M}$ vectors induced by the wave about the direction of propagation. For more general constitutive relations – for example, a chiral medium described by $\mathbf{D}=\epsilon(\mathbf{E}+\beta\boldsymbol{\nabla}\times\mathbf{E})$, $\mathbf{B}=\mu(\mathbf{H}+\beta\boldsymbol{\nabla}\times\mathbf{H}$) – this interpretation is less clear, even when (II) remains a symmetry.
All of these issues stem from the fact that (II) explicitly includes the constitutive parameters $\varepsilon$ and $\mu$ in the transformation. Our approach here is to eschew (II), and to instead consider
@@
as the duality transformation from which to derive the optical helicity inside a general medium. The key to this approach is that the fields $\mathbf{P}$ and $\mathbf{M}$ are not derived from the imposition of constitutive relations, but are explicitly included in an effective model of the material’s internal degrees of freedom (as in [25, 26]).
The explicit inclusion of dynamical material degrees of freedom leads to the natural emergence of a matter component in the medium helicity. We can expect, therefore, the resulting helicity to have, generally, properties different from the case when only the photon degrees of freedom are retained (such as (II)), particularly for excitations that behave in a matter-like rather than photon-like manner. We emphasise that there are no a priori reasons to expect the matter helicity to be closely tied to the circular polarisations of the $\mathbf{A}$, $\mathbf{P}$ or $\mathbf{M}$ fields. Nonetheless, we must also recover the familiar behaviour of the helicity being proportional to the photon-energy density for excitations that are photon-like. The approach taken here is the helicity counterpart of treating the electromagnetic and matter degrees of freedom on equal footing in the energy density, which is required for the optical Thomas-Reiche-Kuhn sum rules [27] to be fulfilled.
III The Helicity Operator as a Generator
In order to enforce both of the transformations in Eq. (II) and, thereby, generate the Heaviside-Larmor transformation (I), we need to extend the form of the helicity and of its flux to include both field and matter components. This suggests that the total helicity density should take the form
$h=\frac{1}{2}\left(\mathbf{A}\cdot\mathbf{B}-\mathbf{C}\cdot\mathbf{D}\right)+\boldsymbol{\Pi}^{P}\cdot\mathbf{M}-\boldsymbol{\Pi}^{M}\cdot\mathbf{P},\\$ (9)
where $\boldsymbol{\Pi}^{P}$ and $\boldsymbol{\Pi}^{M}$ are the canonical momenta conjugate to $\mathbf{P}$ and $\mathbf{M}$ respectively. This helicity now includes an explicit contribution from the matter, as well as from the electromagnetic field. We base this form of the helicity density on the fact its volume integral generates the full Heaviside-Larmor transformation (I). The simplest method to confirm this is to promote our fields to quantum mechanical operators, where
$[\widehat{A}_{i}(\mathbf{x}),\widehat{\Pi}_{j}^{A}(\mathbf{x}^{\prime})]=[\widehat{A}_{i}(\mathbf{x}),-\widehat{D}_{j}(\mathbf{x}^{\prime})]=i\delta_{ij}^{\perp}(\mathbf{x}-\mathbf{x}^{\prime}),$ (10)
$[\widehat{P}_{i}(\mathbf{x}),\widehat{\Pi}_{j}^{P}(\mathbf{x}^{\prime})]=i\delta_{ij}^{\perp}(\mathbf{x}-\mathbf{x}^{\prime}),$ (11)
$[\widehat{M}_{i}(\mathbf{x}),\widehat{\Pi}_{j}^{M}(\mathbf{x}^{\prime})]=i\delta_{ij}^{\perp}(\mathbf{x}-\mathbf{x}^{\prime}).$ (12)
We then find
@@ III The Helicity Operator as a Generator
to first order in $\theta$ and where all fields are transverse. The infinitesimal transformations of $\mathbf{D}$ and $\mathbf{B}$ follow directly.
We have taken all fields to be transverse, as we here wish to study the consequences of this definition of helicity in the simplest possible case, that of a homogeneous medium. However, we note that helicity density in Eq. (9) will always only generate the duality transform for the transverse component of the electromagnetic fields, as $\mathbf{D}$ and $\mathbf{B}$ are necessarily transverse. This is a consequence of the fact that Maxwell’s equations (1) only specify the transverse components of the electromagnetic fields: $\mathbf{D}$ and $\mathbf{B}$ are transverse, and $\mathbf{E}$ and $\mathbf{H}$ appear only through their curls, and so the longitudinal parts do not enter into the dynamics. From the perspective of the helicity as a generator, it is technically possible to include the longitudinal components of the matter parts of the helicity density, i.e. the polarisation $\mathbf{P}$ and magnetisation $\mathbf{M}$ and associated momenta, and promote the respective commutators to full $\delta$-functions. This may be important for the extension to inhomogeneous media. This is, however, outside the scope of what we wish to study here and we leave it to future work.
Finally, we also find that the flux density of our generalised helicity density is simply
@@
which we recognise as the density of the electromagnetic spin. For situations in which the helicity is locally conserved, the helicity density and this flux satisfy the local conservation law
@@
as may be derived directly from Noether’s theorem. The details of this are given in Appendix A.
IV Examples of helicity in media
@@
For simplicity, we here consider here a space filled by a homogeneous lossless magnetodielectric, with a single electric and a single magnetic resonance. The polarisation and magnetisation fields satisfy the equations of motion
@@
where $\omega_{E}$ and $\omega_{M}$ are the resonance frequencies, and $\alpha$ and $\beta$ are constants. In frequency-space, it is clear that these equations imply the constitutive relations $\mathbf{D}=\varepsilon\mathbf{E}$ and $\mathbf{B}=\mu\mathbf{H}$, with
@@
The equations (19), along with Maxwell’s equations, may be obtained from the Lagrangian density
@@
$+\frac{1}{2\alpha^{2}}\big[(\partial_{t}\mathbf{P})^{2}-\omega_{E}^{2}\mathbf{P}^{2}\big]+\frac{1}{2\beta^{2}}\big[(\partial_{t}\mathbf{M})^{2}-(\omega_{M}^{2}+\beta^{2})\mathbf{M}^{2}\big],$
and we take this Lagrangian as a starting point from which to derive an expression for the optical helicity. To begin, we note that the Lagrangian is invariant under the duality transform (II) if (and only if) $\alpha=\beta$ and $\omega_{E}=\omega_{M}$. If we set $\alpha=\beta$ and $\omega_{E}=\omega_{M}=\omega_{0}$, then we can apply Noether’s theorem to recover the helicity directly, as shown in Appendix A. Having established this definition in the dual-symmetric case, we may then apply it in the general case where $\omega_{E}\neq\omega_{M}$ or $\alpha\neq\beta$, and the quantity is no longer locally conserved.
For our chosen model Lagrangian, the canonical momenta corresponding to the polarisation and magnetisation are given by
@@
$\mathbf{\Pi}^{M}\equiv\frac{\partial\mathcal{L}}{\partial\dot{\mathbf{M}}}=\partial_{t}\mathbf{M}/\beta^{2}.$ (23)
These may therefore be interpreted as bound polarisation and magnetisation currents for the Lagrangian (21).
V Quantised helicity
We turn next to a microscopic quantum description of the helicity. In free space, the total helicity can be expressed as the difference between the number of left- and right-handed circularly polarised photons [28, 14, 29],
$\int d^{3}r\,\hat{h}_{\text{free}}=\int d^{3}k\sum_{\lambda=\pm}\lambda\;\hat{a}^{\dagger}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k}),$ (24)
where $\hat{a}_{\lambda}^{\dagger}(\mathbf{k})$ and $\hat{a}_{\lambda}(\mathbf{k})$ are the creation/annihilation operators for a photon of wavevector $\mathbf{k}$, and $\lambda=\pm 1$ labels the two circular polarisations. This is another way of stating that the helicity per photon equals $\pm$ the photon energy divided by its frequency, as the energy density per photon in some mode is simply $k\,\hat{a}^{\dagger}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}(\mathbf{k})$. We will find that the helicity operator defined by equation (9) admits a similar interpretation in a dual-symmetric medium: it is equal to the difference in the number of left- and right-handed polaritons in the system. In other words, a polariton carries helicity that equals $\pm$ the total energy density over the polariton frequency. We shall also see that the relationship is not quite so simple in a general medium, as the matter part of the helicity can contribute with a different sign to the electromagnetic part.
To begin, we diagonalise the system and examine the polariton modes. From the Lagrangian (21), we obtain the Hamiltonian
$+\widetilde{\omega}_{E}\hat{b}^{\dagger}_{\lambda}(\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})+\widetilde{\omega}_{M}\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})$
$-i\frac{\alpha}{2}\sqrt{\frac{k}{\widetilde{\omega}_{E}}}\left[\hat{a}_{\lambda}(\mathbf{k})-\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{b}_{\lambda}(-\mathbf{k})+\hat{b}^{\dagger}_{\lambda}(\mathbf{k})\right]$
$-\lambda\frac{\beta}{2}\sqrt{\frac{k}{\widetilde{\omega}_{M}}}\left[\hat{a}_{\lambda}(\mathbf{k})+\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{c}_{\lambda}(-\mathbf{k})+\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\right].$
where $\hat{a}_{\lambda}(\mathbf{k})$, $\hat{b}_{\lambda}(\mathbf{k})$ and $\hat{c}_{\lambda}(\mathbf{k})$ are the annihilation operators corresponding to the fields $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$ respectively. We have also defined $\widetilde{\omega}_{E}^{2}=\omega_{E}^{2}+\alpha^{2}$ and $\widetilde{\omega}_{M}^{2}=\omega_{M}^{2}+\beta^{2}$. We define the circular polarisation vectors $\mathbf{e}_{\pm}$ with the convention that $\mathbf{e}_{\lambda}(\mathbf{k})\cdot\mathbf{e}_{\lambda^{\prime}}^{*}(\mathbf{k})=\delta_{\lambda\lambda^{\prime}}$, $\mathbf{e}_{\lambda}(\mathbf{k})\cdot\mathbf{k}=0$ for all $\lambda$, and $\mathbf{e}_{\lambda}(-\mathbf{k})=-\mathbf{e}_{-\lambda}(\mathbf{k})$. Our continuum annihilation and creation operators satisfy the bosonic commutation relations $\hat{a}_{\lambda}(\mathbf{k}),\hat{a}^{\dagger}_{\lambda^{\prime}}(\mathbf{k}^{\prime})]=\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, with similar relations for $\hat{b}_{\lambda}(\mathbf{k})$ and $\hat{c}_{\lambda}(\mathbf{k})$, and all other commutators equal to zero. The explicit expansions of $\hat{\mathbf{A}}$, $\hat{\mathbf{P}}$ and $\hat{\mathbf{M}}$ used to obtain (25) from (21) are given in Appendix B.
We consider first the helicity in a dual-symmetric medium, and show that the total helicity is naturally expressible in terms of the polariton modes of the system. As mentioned above, in a dual-symmetric medium, the helicity operator defined by equation (9) is simply equal to the difference between the number of left- and right-handed polaritons in the system. We will then go on to show how this picture is modified in more realistic media.
V.1 Helicity in dual symmetric media
In order to specialise (25) to a dual-symmetric medium, we consider the case where $\omega_{E}=\omega_{M}\equiv\omega_{0}$ and $\alpha=\beta$. The Hamiltonian then becomes